OpenStudy (anonymous):

Assets: $ 43,000 Liability: ? Stock: $6,000 Net Income (loss): $9650 Cash Dividends: $3500 This is how to calculate it... Assets = Liabilities + Equity Equity = Stock + Net Income - Dividends Please show work. I keep getting a negative and a wrong answer. Please help

OpenStudy (anonymous):

Equity = 6000+9650-3500 = $12,150 Assets = liabilities + equity or, 43,000 = liabilities + 12,150 or, Liabilities = 43,000 - 12,150 [by changing the side] or, liabilities = $ 30,850 So simple, you probably made error in changing sign in time of changing the side.